• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Puckett

    Log entry time 04:45:08 on April19,2007

    Entry number 199550

    keyword=tools and BCM screengrab

    Now that the dust has settled, we are running at about 73 muA, which is as high as we can go for a reasonable flow rate and heat load on the 4He cryotarget. Elena is checking the coin. time spectra. We are taking data, and as soon as we can verify the coincidence time we will keep running until we have the necessary data.



    FIGURE 1

    FIGURE 2